OpenStudy (anonymous):

Can someone please show me how to do this? im confused and thank you :)write an equation of a line that passes through the given point and is parallel to graph of the given equation. (2,-1);y=5x-2

jigglypuff314 (jigglypuff314):

y = mx + b is the form for slope-intercept form m = slope and b = y-intercept the line you are looking for would have to have the same slope as the given equation in order for the two lines to be parallel so m = 5 from y = 5x - 2 then make an equation given that slope = m = 5 and the point (2, -1)

OpenStudy (anonymous):

oh , ok I I have this one (0,-5);y=9x so the slope would be 9?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

but how would I find the line that passes through? that's what its asking and idk what it is :/

jigglypuff314 (jigglypuff314):

I'm not quite sure what you are asking, may I ask what the full question was?

OpenStudy (anonymous):

yeah. write an equation of a line that passes through the given point and is parallel to graph of the given equation. this is what its asking and it give me equations

jigglypuff314 (jigglypuff314):

okay, so for given the point (0,-5); and the equation y=9x a line parallel to y = 9x would have to have a slope of 9 then y = 9x + b and temporarily plug in the points (0, -5) into y = 9x + b to find b. so (-5) = 9(0) + b then b = -5 so the equation that they want would be y = 9x - 5

OpenStudy (anonymous):

Then how would I be able to do this one (2,-1);y=5x-2

OpenStudy (anonymous):

would -2 replace B?

jigglypuff314 (jigglypuff314):

from y=5x-2, the slope would be 5 then with y = 5x + b you temporarily plug in the points (2, -1) into y = 5x + b to find b so (-1) = 5(2) + b but b would Not be -2

OpenStudy (anonymous):

so the equation would be -1=10+b? sorry if i'm askig too much question

jigglypuff314 (jigglypuff314):

you would solve for b from -1 = 10 + b then plug in what you got from b back into the y = 5x + b

OpenStudy (anonymous):

oh ok thank you so much. May I ask you on more quesio?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

ok, so this one asks us graph the equation but the equations are like y+3=2(x-1)

OpenStudy (anonymous):

can show me example so I can do the rest by myself :)

jigglypuff314 (jigglypuff314):

ah yes, okay, so that would be in point-slope form: y - y1 = m(x - x1) when m = slope and given point on line (x1, y1) so for an equation in point-slope form like y + 3 = 2(x - 1) you would graph it knowing that the slope = 2 and a point is (1, -3)

OpenStudy (anonymous):

cool thank you again.

jigglypuff314 (jigglypuff314):

you're welcome :)